WHAT WILL BE THE CONCENTRATION OF CHLORIDE ION IF 200 ml of 0.2 M CaCl2 IS MIXED  WITH 300 ml OF 0.1 M  NaCl SOLUTION ?

Grade:12

1 Answers

mrunali
6 years ago
HERE IS THE DIRECT FORMULA TO CALCULATE THE RESULTANT CONCENTRATION-
M(V1 + V2) = M1V1 + M2V2
HERE M1 IS THE CONCENTRATION OF CHLORIDE FROM CaCl2 = 2 * 0.2 (as in CaCl2 2 Cl are present ) and V1=200
M2 IS THE CONCENTRATION OF CHLORIDE FROM Nacl = 0.1 AND V2 =300
SO AFTER PUTTING VALUES THE ANSWER WILL BE 
M* 500 = 200*0.4 + 300*0.1
M=110/500
M= 0.22